OpenStudy (anonymous):

What is the solution set for |3x + 12| + 5 = 23

OpenStudy (anonymous):

I think it would be x=2 and x=10

OpenStudy (anonymous):

\[\left| 3x+12 \right|=23\] so you have two solutionss \[3x+12=18\] \[3x=6\] \[x=2\] AND when its negetive \[3x+12=-18\] \[3x=-30\] \[x=-10\]

OpenStudy (anonymous):

THANK YOU BOTH!!!! :)

OpenStudy (anonymous):

Jonask will you assist me with my problem please?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

please click on the problem i posted